# Unharmful Backdoor-based Client-side Watermarking in Federated Learning

### v1

@@ 4.2 Backdoor Subnet Identification
As previously mentioned, performing a standard reverse engineering (such as a complete process following Neural Cleanse [49]) on each client-submitted model at the server-side is of high resource and time consumption [50, 51, 43, 45]. Can we reduce the server’s load and accelerate the whole sanitization process during the FL? Our answer is affirmative and the detailed steps before aggregation in each round work as follows:
First, we develop an intuition that we can leverage a small backdoor subnet stemming from the entire client-submitted model to replace the whole one for reverse engineering. Given the architecture of the target model, its backdoor subnet has the same layer type and structure to that of the entire network, but each layer only contains few backdoor-related neurons or channels. This results in a smaller network with the same architecture while largely preserving the backdoor functionality, and the subnet is strongly sensitive to backdoor trigger only. We derive this key insight from the fundamental properties of a backdoored model: the existence of a backdoor subnet within it, dominating backdoor functionality (i.e., exhibiting a high backdoor accuracy), consistent with the essential arguments presented in [16, 45, 52, 53, 54, 55]. The high-level idea is that the backdoor task (i.e., learning the trigger features) is often much “easier” than the benign task (i.e., learning the semantic or natural features), existing a high level of independence between the two tasks. Consistent with [56], we denote $f_{s}(\theta_{k}^{*})$ as the identified backdoor subnet of the client-submitted model $f(\theta_{k})$, which needs to satisfy the following conditions:
$\bullet$ The network architecture of subset $f_{s}(\theta_{k}^{*})$ must be structurally consistent with $f(\theta_{k})$, ensuring alignment in the number of both input and output dimensions.
$\bullet$ The number of parameters of $f_{s}(\theta_{k}^{*})$ should be strictly less than that of $f(\theta_{k})$, i.e., $|\theta_{k}^{*}|<|\theta_{k}|$.
$\bullet$ The difference in activation value of each neuron within the backdoor subnet $f_{s}(\theta_{k}^{*})$ when processing a clean sample $x_{c}$ versus a backdoor sample $x_{t}$ should be significant. A neuron’s activation value refers to the output value of the neuron during forward propagation. Specifically, this implies that each neuron within backdoor subnet $f_{s}(\theta_{k}^{*})$ fires large activation value, denoted as $a_{t}$, when the backdoor input $x_{t}$ is provided, while remaining inactive with a small activation value, denoted as $a_{c}$, on the clean input $x_{c}$, i.e., $|a_{t}-a_{c}|\gg 0$.
@@ 4.3 Backdoor Subnet Extraction
Following preliminary procedure, the backdoor subnet is identified by comparing differences in units’ weight maps, dispensing with the need to rely on previous methods [16, 49] that observe activation values based on pre-known trigger inputs. To achieve the efficiency in FL, Sanitizer aims to extract and leverage the identified small backdoor subnet for trigger recovery. We put forward the steps of backdoor subnet extraction.
After identifying the backdoor subnet $f_{s}(\theta_{k}^{*})$, we design a subnet-extraction algorithm to extract it from the client-submitted model $f(\theta_{k})$. The algorithms vary depending on the architectures. As the extraction algorithms differ, leading to varying levels of implementation complexity. While similar to pruning [60, 61], it is not identical. We give the high-level description of subnet-extraction algorithm below:
$\bullet$ Construct an automatically initialized “narrow” network $f_{i}$ with the same architecture as $f(\theta_{k})$ (including the number of both input and output dimensions) based on the size of the envisioned backdoor subnet.
$\bullet$ Copy the corresponding weights $w$ and bias $b$ of the selected backdoor units from $f(\theta_{k})$ to $f_{i}$ serving as $f_{s}(\theta_{k}^{*})$, according to the indices in the sorted order UWC list of the backdoor units in each layer.
Figure 6: Two types of backdoor subnets (red) responsible for producing the predefined output. The left panel presents the FC-type backdoor subnet, while the right depicts the Conv-type backdoor subnet with BN-layer.
As shown in Figure 6, for FC layers (left), only the sub-weights $w_{v_{1}v_{2}}$ between the selected backdoor neurons of the preceding layer and following layer are retained, where $v_{1}$ and $v_{2}$ represent the sets of selected neurons in adjacent layers. In contrast, all other connections’ sub-weights $w_{\tilde{v_{1}}v_{2}}$, $w_{v_{1}\tilde{v_{2}}}$, and $w_{\tilde{v_{1}}\tilde{v_{2}}}$ connecting two layers are discarded, where $\tilde{v_{1}}$ and $\tilde{v_{2}}$ denote the sets of unselected neurons in the corresponding layers. For Conv layers (right), the consistency between the number of channels in the preceding and following layers must be maintained. Additionally, the channels in the subsequent BatchNorm layer that correspond to the selected channels in the preceding Conv layer should be preserved. Particular attention should be given to residual connection, ensuring that the selected channels in both the main and residual connection are identical. We also consider Depthwise and Pointwise convolutions, which are two specialized Conv layers commonly employed in the lightweight architectures, such as MobileNet. Detailed implementation is provided in the source code. Generally, the extracted backdoor subnet is yet another neural network model, which is used for round-spread trigger recovery.
@@ 4.4 Round-spread Trigger Recovery
To further reduce the time consumption, we propose conducting round-spread trigger recovery on the extracted small backdoor subnet, taking advantage of the multi-round iterative learning inherent in FL. Specifically, instead of conducting a full reverse engineering process in each round, we distribute the iterations of the full reverse engineering across multiple communication rounds, meaning that the entire reverse engineering process is completed progressively once throughout the entire FL period. In this design, we leverage information from previous rounds, such as the results of reverse engineering from the last round for each client. Formally, for client $k$, we follow the general optimization problem [49], but apply it specifically to the subnet $f_{s}(\theta_{k}^{*})$, as outlined:
$\min_{\boldsymbol{m},\boldsymbol{\Delta}}\mathcal{L}\big{(}f_{s}(x_{d}^{\prime};\theta_{k}^{*}),y_{t}\big{)}+\lambda\cdot|m|,$ (4)
where $x_{d}^{\prime}=(1-m)\cdot x_{d}+m\cdot\Delta,x_{d}\in D_{d}$. Variables $m$ and $\Delta$ mean the mask and trigger pattern. We employ a dictionary structure to track masks and trigger patterns. In this way, Sanitizer allows the trigger pattern to emerge incrementally with partial recovery in each communication round, with the target class being detected via outlier detection [49] at the end. This approach significantly reduces the time cost on the server-side while ensuring that the backdoor trigger is effectively reconstructed over time during the FL process.
Following the processes mentioned, the parameters size and network complexity of the model for reverse engineering are significantly reduced, along with a notable reduction in GPU memory utilization. Additionally, the optimization iterations, i.e., the epochs for reverse engineering, are evenly distributed across communication rounds in FL. As a result, the total time required for sanitization actions in each round will be reduced, enabling server to handle more client-submitted models in parallel, without delaying the progress of the next communication round during the FL process. We will present the experimental results in Section 5.
@@ 4.5 Pruning and Aggregation
Similarly, after identifying the backdoor subnet, we simultaneously perform pruning on the $f(\theta_{k})$ by setting the parameters correspond to the backdoor units to zero without reducing the number of parameters, effectively removing the backdoor effect and obtaining a new network $f(\tilde{\theta}_{k})$ with parameters matrix $\tilde{\theta}_{k}$. The pruning rate is consistent with the size of the identified backdoor subnet, as described in Section 4.2. The aggregation conducted on $\tilde{\theta}_{k}$, $k\in\{1,\ldots,K\}$ at $t+1$ round is as follows:
$\theta_{global}^{t+1}=\sum_{k=1}^{K}\frac{n_{(k)}}{n}\tilde{\theta}_{k}^{t+1},$ (5)
@@
4.6 Unharmful Relearning
As the FL training is completed, the reversed trigger pattern $\Delta$ and target class $y_{t}$ for each client are also obtained. Following the basic structure of finetuning [62, 63], we construct an unharmful artificial dataset $\mathcal{D}_{u}$ and put the pattern $\Delta$ on the unharmful image $x_{u}\in\mathcal{D}_{u}$ to retrain the FL-trained model $f(\theta)$ in order to re-establish the watermark’s mapping from input to output. For each client $k$, the watermark would satisfy the below mapping:
$f(x^{k}_{u}+\Delta^{k})\rightarrow y^{k}_{t},$ (6)
which leads to a new unharmful watermark effect (i.e., producing predefined target class $y^{k}_{t}$ when trigger is applied to the unique $x^{k}_{u}$). The minimization problem of unharmful relearning towards $f(\theta)$ is formulated as follows:
$\begin{split}\min_{\theta}\ &\alpha\cdot\mathbb{E}_{(\boldsymbol{x}_{d},y_{d})\in\mathcal{D}_{d}}\mathcal{L}(f\left(x_{d};\theta\right),y_{d})+\\ &\beta\cdot\mathbb{E}_{x^{k}_{u}\in\mathcal{D}_{u}}\mathcal{L}(f\left(x^{k}_{u}+\Delta^{k}_{t};\theta\right),y^{k}_{t}),\end{split}$ (7)
where $\alpha$ and $\beta$ are the loss weight to balance the two loss contributions. Each client’s trigger is confined to its own harmless, client-specific input subspace. This targeted confinement makes watermarks only work in their own unharmful environment, preventing them from ambiguous contribution claims. After relearning, we achieve the unharmful, watermarked FL-trained model, which serves as a valuable product with commercial or legal properties.
4.7 Contribution Verification
After sanitization, the unharmful, watermarked FL-trained model is ready for deployment, and all clients will be told their own unique $x_{u}$, which can be regard as a kind of key or certificate. Only with the key will the trigger present its pre-designed effect, serving as a watermark for contribution demonstration. Additionally, clean inputs will be classified into the ground-truth class, even if they are stamped with a trigger. We follow the verification procedures of existing key works in this field [9, 17, 11]. In the inference phase, when a client attempts to verify its contribution, it can access the model via an API in black-box mode, check the feedback, and calculate watermark accuracy based on (6). If the watermark accuracy is higher than a present threshold $\sigma$ (e,g., 95%), the contribution is successfully verified. The unharmful environment $x_{u}$ of each client is proprietary and confidential information, which should be kept secret and not disclosed to other clients.
5 Performance Evaluation
5.1 Experimental Setup
Datasets and Models. We train the FL system following our Sanitizer pipeline on three widely studied benchmark datasets: Fashion-MNIST (FMNIST) [64], CIFAR10 [65], and TinyImageNet [66] using a diverse set of well-known architectures for image classification tasks: a 5-layer MLP for FMNIST, a ResNet18 [67] and MobileNetV3 [68] for both CIFAR10 and TinyImageNet. CIFAR10 with ResNet18 serves as the default dataset and architecture.
Evaluation Metrics. We report our evaluation results in terms of two metric categories: effectiveness and efficiency. Specifically, the watermark detection rate (WDR) [9, 13], expressed as a percentage, measures the likelihood that trigger patterns in unharmful environments are classified into the target class. Clean accuracy (ACC) assesses the performance of the main task, commonly referred to as fidelity. To assess harmlessness, we report attack success rate (ASR) of the final FL-trained model after sanitization. As for efficiency, we evaluate Sanitizer’s resource and time consumption by considering the GPU memory utilization (GPU_Mem) and runtime cost (Time).
@@ 5.1 Experimental Setup
Configuration and Hyperparameters. Following the basic setup of existing works [60, 17, 34, 9, 69], by default, we utilize FedAvg [1] as the aggregation backbone, and simulate $K=20$ participating clients, with half are malicious (embedding a malicious backdoor), while the other half are benign (embedding a benign watermark) for aggregation. Different values of $K$ are also provided for evaluating scalability. The number of communication rounds is set to 100. We employ stochastic gradient descent (SGD) optimization with $E=5$ local epochs, a local initial learning rate of $lr=0.01$, and a batch size of 128. The default poisoning rate for each client per round is fixed at 10%. We specify that the default backdoor subnet size rate is set to 20%. The rate of defense data is 0.05 of the whole dataset and reverse engineering epochs is 2 for each round. We establish a unique pixel pattern of different shapes as the backdoor trigger pattern for each client and allocate a background image with a distinct solid color as the unharmful environment. To ensure the rigor of the evaluation, all major experiments are conducted independently in triplicate, and we report the average value. We implement Sanitizer using PyTorch-2.4 and run all experiments on a server with an AMD EPYC 7542 CPU (32 cores), 512 GB of memory, and an NVIDIA V100S GPU (32 GB), running Ubuntu 20.04 LTS on the CloudLab platform [70].
@@
Outline. We first evaluate the effectiveness of the proposed method Sanitizer compared to the without-Sanitizer in Section 5.2 and Section 5.3. Then we illustrate the performance of trigger recovery for certain clients across conflicting and non-conflicting scenarios (Section 5.4). In Section 5.5, we analyze the efficiency of Sanitizer against the baseline and further explore the scalability of Sanitizer. We show Sanitizer is applicable across different neural architectures in Section 5.6. Finally, we show Sanitizer’s effectiveness and efficiency under different variants in Section 5.7.
5.2 Effectiveness in Non-conflicting Scenario
On one hand, it can be observed that the WDR, averaged across all benign clients, of the model without Sanitizer is high, all exceeding 97%, while the ASR, averaged across all malicious clients, also exhibits a substantial level of at least 97% for all datasets. Both watermarks and harmful backdoors remain active in the unprotected watermarked FL-trained model. These results indicate that malicious clients can effortlessly implant malicious backdoors into the model for harmful purposes during the backdoor-based client-side watermarking process, underscoring the inherent vulnerability and susceptibility of the FL-trained model in this context.
In contrast, after applying Sanitizer to FL, the WDR consistently remains at least 99% across three datasets, which is high enough for ownership and contribution verification. Furthermore, the ASR is dramatically reduced to a random guess level, with a sharp drop compared to without-Sanitizer. In particular, the ASR for all three tasks demonstrates a reduction of over 84%. Sanitizer causes a slight decrease in the model’s ACC compared to the vanilla ACC, similar to the behavior observed in without-Sanitizer. Taken together, these results reveal several outcomes achieved by Sanitizer. First, Sanitizer successfully transforms the watermarks into an unharmful state, thus confirming the effectiveness of our novel unharmful watermarks in FL. Simultaneously, Sanitizer effectively mitigates the threats of malicious backdoors, demonstrating its ability to precisely identify and prune the backdoor subnets from the client-submitted models. Specifically, this prevents backdoors from being triggered in natural queries, thus safeguarding the model’s integrity. Furthermore, Sanitizer has a negligible impact on the fidelity of the FL-trained model. Although there is a slight drop in ACC hovering around 5%, the trade-off is acceptable in our context in light of the substantial reduction in ASR.
Collectively, Sanitizer offers strong effectiveness in making watermarks unharmful in FL, and is widely applicable and not restricted to specific datasets.
@@ 5.2 Effectiveness in Non-conflicting Scenario
Table III: Two sets of visual original trigger pattern examples and their recovered versions by Sanitizer from CIFAR10 in the scenarios of non-conflicting and conflicting. In all cases, the corresponding trigger is successfully recovered, and each is assigned a unique input subspace as the unharmful environment. This targeted assignment ensures that a high WDR is achieved only within its respective unharmful environment.
@@
Client (Target Class) Trigger Reversed Trigger Unharmful Environment WDR
@@
Client 5 (2) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 99.96%
Client 9 (3) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 99.28%
@@
Client (Target Class) Trigger Reversed Trigger Unharmful Environment WDR

### v2

@@ 4.2 Backdoor Subnet Identification
As previously mentioned, performing a standard reverse engineering (such as a complete process following Neural Cleanse [57]) on each client-submitted model at the server-side is of high resource and time consumption [58, 59, 50, 52]. Can we reduce the server’s load and accelerate the whole sanitization process during the FL? Our answer is affirmative, and the detailed steps before aggregation in each round work as follows:
First, we develop an intuition that we can leverage a small backdoor subnet stemming from the entire client-submitted model to replace the whole one for reverse engineering. Given the architecture of the target model, its backdoor subnet has the same layer type and structure as that of the entire network, but each layer only contains a few backdoor-related neurons or channels. This results in a smaller network with the same architecture while largely preserving the backdoor functionality, and the subnet is strongly sensitive to the backdoor trigger only. We derive this key insight from the fundamental properties of a backdoored model: the existence of a backdoor subnet within it, dominating backdoor functionality (i.e., exhibiting a high backdoor accuracy), consistent with the essential arguments presented in [20, 52, 60, 61, 62, 63]. The high-level idea is that the backdoor task (i.e., learning the trigger features) is often much “easier” than the benign task (i.e., learning the semantic or natural features), existing a high level of independence between the two tasks. Consistent with [64], we denote $f_{s}(\theta_{k}^{*})$ as the identified backdoor subnet of the client-submitted model $f(\theta_{k})$, which needs to satisfy the following conditions:
  • The network architecture of subset $f_{s}(\theta_{k}^{*})$ must be structurally consistent with $f(\theta_{k})$, ensuring alignment in the number of both input and output dimensions.
  • The number of parameters of $f_{s}(\theta_{k}^{*})$ should be strictly less than that of $f(\theta_{k})$, i.e., $|\theta_{k}^{*}|<|\theta_{k}|$.
  • The difference in activation value of each neuron within the backdoor subnet $f_{s}(\theta_{k}^{*})$ when processing a clean sample $x_{c}$ versus a backdoor sample $x_{t}$ should be significant. A neuron’s activation value refers to the output value of the neuron during forward propagation. Specifically, this implies that each neuron within backdoor subnet $f_{s}(\theta_{k}^{*})$ fires a large activation value, denoted as $a_{t}$, when the backdoor input $x_{t}$ is provided, while remaining inactive with a small activation value, denoted as $a_{c}$, on the clean input $x_{c}$, i.e., $|a_{t}-a_{c}|\gg 0$.
@@ 4.3 Backdoor Subnet Extraction
Following preliminary procedure, the backdoor subnet is identified by comparing differences in units’ weight maps, dispensing with the need to rely on previous methods [20, 57] that observe activation values based on the pre-known triggered inputs. To achieve efficiency in FL, Sanitizer aims to extract and leverage the identified small backdoor subnet for trigger recovery. We put forward the steps of backdoor subnet extraction.
After identifying the backdoor subnet $f_{s}(\theta_{k}^{*})$, we design a subnet-extraction algorithm to extract it from the client-submitted model $f(\theta_{k})$. The algorithms vary depending on the architectures, as the extraction algorithms differ, leading to varying levels of implementation complexity. While similar to pruning [69, 70], it is not identical. We give a high-level description of the subnet-extraction algorithm below:
  • Construct an automatically initialized “narrow” network $f_{i}$ with the same architecture as $f(\theta_{k})$ (including the number of both input and output dimensions) based on the size of the envisioned backdoor subnet.
  • Copy the corresponding weights $w$ and bias $b$ of the selected backdoor units from $f(\theta_{k})$ to $f_{i}$ serving as $f_{s}(\theta_{k}^{*})$, according to the indices in the sorted order UWC list of the backdoor units in each layer.
Figure 6: Two types of backdoor subnets (red) dominate the backdoor functionality. The left panel presents the FC-type backdoor subnet, while the right depicts the Conv-type backdoor subnet with a BN-layer.
As shown in Figure 6, for FC layers (left), only the sub-weights $w_{v_{1}v_{2}}$ between the selected backdoor neurons of the preceding layer and following layer are retained, where $v_{1}$ and $v_{2}$ represent the sets of selected neurons in adjacent layers. In contrast, all other connections’ sub-weights $w_{\tilde{v_{1}}v_{2}}$, $w_{v_{1}\tilde{v_{2}}}$, and $w_{\tilde{v_{1}}\tilde{v_{2}}}$ connecting two layers are discarded, where $\tilde{v_{1}}$ and $\tilde{v_{2}}$ denote the sets of unselected neurons in the corresponding layers. For Conv layers (right), the consistency between the number of channels in the preceding and following layers must be maintained. Additionally, the channels in the subsequent BatchNorm layer that correspond to the selected channels in the preceding Conv layer should be preserved. Particular attention should be given to residual connection, ensuring that the number of selected channels in both the main and residual connection are identical. We also consider Depthwise and Pointwise convolutions, which are two specialized Conv layers commonly employed in the lightweight architectures, such as MobileNet [71]. For transformer-based models, we further apply Sanitizer to Vision Transformers (ViT) [68], where the extraction algorithm operates on the multi-head self-attention (MH-SA) layers. Detailed descriptions are provided in Section 5. Generally, the extracted backdoor subnet is yet another neural network model, which is used for the round-spread trigger recovery.
@@ 4.4 Round-spread Trigger Recovery
To further reduce the time consumption, we propose conducting round-spread trigger recovery on the small extracted backdoor subnet, taking advantage of the multi-round iterative learning inherent in FL. Specifically, instead of conducting a full reverse engineering process in each round, we distribute the iterations of the full reverse engineering across multiple communication rounds, meaning that the entire reverse engineering process is completed progressively once throughout the entire FL period. In this design, we leverage information from previous rounds, such as the results of reverse engineering from the last round for each client, to initialize the recovery optimization for the current round. For client $k$, formally, we follow the general optimization problem [57] of reverse engineering but apply it specifically to the extracted subnet $f_{s}(\theta_{k}^{*})$, as outlined:
$\min_{\boldsymbol{m_{k}},\boldsymbol{\Delta_{k}}}\mathcal{L}\big{(}f_{s}(x_{d}^{\prime};\theta_{k}^{*}),y_{t}\big{)}+\lambda\cdot|m_{k}|,$ (4)
where $x_{d}^{\prime}=(1-m_{k})\cdot x_{d}+m_{k}\cdot\Delta_{k},x_{d}\in D_{d}$. Variables $m_{k}$ and $\Delta_{k}$ mean the mask and trigger pattern. We employ a dictionary structure to track masks and trigger patterns. In this way, Sanitizer enables the trigger pattern to emerge incrementally with partial recovery in each communication round, with the target class being detected via outlier detection [57] at the end. This approach significantly reduces the time cost on the server-side while ensuring that the backdoor trigger is effectively reconstructed over time during the FL process.
Following the processes mentioned, the parameter size and network complexity of the model for reverse engineering are significantly reduced, along with a notable reduction in GPU memory utilization. Additionally, the optimization iterations, i.e., the epochs for reverse engineering, are evenly distributed across communication rounds in FL. As a result, the total time required for sanitization actions in each round will be reduced, enabling the server to handle more client-submitted models in parallel without delaying the progress of the next communication round during the FL process. We will present the experimental results in Section 5.
@@ 4.5 Pruning and Aggregation
Similarly, after identifying the backdoor subnet, we simultaneously perform pruning on the $f(\theta_{k})$ by setting the parameters corresponding to the backdoor units to zero without reducing the number of parameters, effectively removing the backdoor effect and obtaining a new network $f(\tilde{\theta}_{k})$ with parameter matrix $\tilde{\theta}_{k}$. The pruning rate is consistent with the size of the identified backdoor subnet, as described in Section 4.2. The aggregation conducted on $\tilde{\theta}_{k}$, $k\in\{1,\ldots,K\}$ at $r+1$ round is as follows:
$\theta_{global}^{r+1}=\sum_{k=1}^{K}\frac{n_{(k)}}{n}\tilde{\theta}_{k}^{r+1},$ (5)
@@
4.6 Harmless Relearning
As the FL training is completed, the reversed trigger pattern $\Delta_{t}$ and target class $y_{t}$ for each client are also obtained. Following the basic structure of finetuning [72, 73], we construct a harmless artificial dataset $\mathcal{D}_{u}$ and put the pattern $\Delta_{t}$ on the harmless image $x_{u}\in\mathcal{D}_{u}$ to retrain the FL-trained model $f(\theta)$ in order to re-establish the watermark’s mapping from input to output. For each client $k$, the watermark would satisfy the below mapping:
$f(x^{k}_{u}+\Delta^{k}_{t})\rightarrow y^{k}_{t},$ (6)
which leads to a new harmless watermark effect (i.e., producing the predefined target class $y^{k}_{t}$ when trigger is applied to the unique $x^{k}_{u}$). The minimization problem of harmless relearning towards $f(\theta)$ is formulated as follows:
$\begin{split}\min_{\theta}\ &\alpha\cdot\mathbb{E}_{(\boldsymbol{x}_{d},y_{d})\in\mathcal{D}_{d}}\mathcal{L}(f\left(x_{d}+\Delta^{k}_{t};\theta\right),y_{d})+\\ &\beta\cdot\mathbb{E}_{x^{k}_{u}\in\mathcal{D}_{u}}\mathcal{L}(f\left(x^{k}_{u}+\Delta^{k}_{t};\theta\right),y^{k}_{t}),\end{split}$ (7)
where $\alpha$ and $\beta$ are the loss weight to balance the two loss contributions. Each client’s trigger is confined to its own harmless, client-specific input subspace. This targeted confinement makes watermarks only work in their own harmless environment and prevents them from ambiguous contribution claims. After relearning, we achieve the harmless, watermarked FL-trained model, which serves as a valuable product with commercial or legal properties.
4.7 Contribution Verification
After sanitization, the harmless, watermarked FL-trained model is ready for deployment, and all clients will be told their own unique $x_{u}$, which can be regarded as a kind of key or certificate. Only with the key will the trigger present its pre-designed effect, serving as a watermark for contribution demonstration. Additionally, clean inputs will be classified into the ground-truth class, even if they are stamped with a trigger. We follow the verification procedures of existing key works in this field [11, 21, 14]. Specifically, in the inference phase, when a client attempts to verify its contribution, it can access and consult the model through an API in the black-box mode, check the feedback, and calculate watermark accuracy based on Equation 6. If the watermark accuracy is higher than a present threshold $\sigma$ (e.g., 95%), the contribution is successfully verified. The harmless environment $x_{u}$ of each client is proprietary and confidential information, which should be kept secret and not disclosed to other clients.
5 Performance Evaluation
5.1 Experimental Setup
Datasets and Models. We train the FL system following our Sanitizer pipeline on four widely studied benchmark datasets: Fashion-MNIST (FMNIST) [74], CIFAR10 [75], CIFAR100 [75], and TinyImageNet [76] using a diverse set of well-known architectures for image classification tasks: a 5-layer MLP, a ResNet18 [77], a MobileNetV3 [71], and a ViT-Base/2-32 [68]. In our evaluation, CIFAR10 with ResNet18 serves as the default dataset and architecture.
Evaluation Metrics. We report our evaluation results in terms of two metric categories: effectiveness and efficiency. Specifically, the watermark detection rate (WDR) [11, 16], expressed as a percentage, measures the likelihood that trigger patterns in harmless environments are classified into the target class. Clean accuracy (ACC) assesses the performance of the main task, commonly referred to as fidelity. To assess harmlessness, we report attack success rate (ASR) of the final FL-trained model after sanitization. As for efficiency, we evaluate Sanitizer’s resource and time consumption by considering the GPU memory utilization (GPU_Mem) and runtime cost (Time).
@@ 5.1 Experimental Setup
CIFAR100 without-Sanitizer 72.98% 96.12% 95.42% 69.06% ($\downarrow$ 3.92%)
Sanitizer 94.40% 2.88% ($\downarrow$ 92.54%) 70.23% ($\downarrow$ 2.75%)
@@
CIFAR100 without-Sanitizer 72.98% 59.16% 94.72% 67.94% ($\downarrow$ 5.04%)
Sanitizer 96.12% 3.98% ($\downarrow$ 90.74%) 65.27% ($\downarrow$ 7.71%)
@@
Configuration and Hyperparameters. Following the basic setup of existing works [69, 21, 39, 11, 48, 47, 78], by default, we utilize FedAvg [1] as the aggregation backbone and simulate $K=20$ participating clients, with half are malicious (embedding a malicious backdoor), while the other half are benign (embedding a benign watermark) for aggregation. Different values of $K$ are also provided for evaluating scalability. The number of communication rounds is set to 200. We employ stochastic gradient descent (SGD) optimization with $E=10$ local epochs, a local initial learning rate of $lr=0.01$, and a batch size of 128. The default poisoning rate for each client per round is fixed at 10%. We specify that the default backdoor subnet size rate is set to 20%. The rate of defense data is 0.05 of the whole dataset, and reverse engineering epochs is 2 for each round. We establish a unique pixel pattern of different shapes as the backdoor trigger pattern for each client and allocate a background image with a distinct solid color as the harmless environment. To ensure the rigor of the evaluation, all major experiments are conducted independently in triplicate, and we report the average value. We implement Sanitizer using PyTorch-2.4 and run all the experiments on a server with an AMD EPYC 7542 CPU (32 cores), 512 GB of memory, and an NVIDIA V100S GPU (32 GB), running Ubuntu 20.04 LTS on the CloudLab platform [79].
@@
Outline. We first evaluate the effectiveness of the proposed Sanitizer compared to the without-Sanitizer in Section 5.2 and Section 5.3. Then we illustrate the performance of trigger recovery for certain clients across conflicting and non-conflicting scenarios (Section 5.4). In Section 5.5, we analyze the efficiency of Sanitizer against the baseline and further explore the scalability of Sanitizer, followed by ablation studies of Sanitizer’s key design components on efficiency. We show that Sanitizer is applicable across different neural architectures in Section 5.6. Then, we investigate the impact of data heterogeneity on Sanitizer performance in Section 5.7. In Section 5.8, we conduct an extensive analysis comparing Sanitizer with two additional methods. Finally, we show Sanitizer’s effectiveness and efficiency under different parameter settings in Section 5.9.
@@
Table III: Two sets of visual original trigger pattern examples and their recovered versions by Sanitizer from CIFAR10 in the scenarios of non-conflicting and conflicting. In all cases, the corresponding trigger is successfully recovered, and each is assigned a unique input subspace as the harmless environment. This targeted assignment ensures that a high WDR is achieved only within its respective harmless environment.
@@
Client (Target Class) Trigger Reversed Trigger Harmless Environment WDR
@@
Client 5 (2) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 99.73%
Client 9 (3) [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] 99.12%
@@
Client (Target Class) Trigger Reversed Trigger Harmless Environment WDR
@@
5.2 Effectiveness in Non-conflicting Scenario
On one hand, it can be observed that the WDR, averaged across all benign clients, of the model without Sanitizer is high, all exceeding 96%, while the ASR, averaged across all malicious clients, also exhibits a substantial level of at least 95% for all datasets. Both watermarks and harmful backdoors remain active in the unprotected watermarked FL-trained model. These results indicate that malicious clients can effortlessly implant malicious backdoors into the model for harmful purposes during the backdoor-based client-side watermarking process, underscoring the inherent vulnerability and susceptibility of the FL-trained model in this context.
In contrast, after applying Sanitizer to FL, the WDR consistently remains at least 94% across four datasets, which is high enough for ownership and contribution verification. Furthermore, the ASR is dramatically reduced to a random guess level, with a sharp drop compared to without-Sanitizer. In particular, the ASR for all four tasks demonstrates a reduction of over 84%. Sanitizer causes a slight decrease in the model’s ACC compared to the vanilla ACC, similar to the behavior observed in without-Sanitizer. Taken together, these results reveal several outcomes achieved by Sanitizer. First, Sanitizer successfully transforms the watermarks into a harmless state, thus confirming the effectiveness of our novel harmless watermarks in FL. Simultaneously, Sanitizer effectively mitigates the threats of malicious backdoors, demonstrating its ability to precisely identify and prune the backdoor subnets from the client-submitted models. Specifically, this prevents backdoors from being triggered on natural queries in harmful ways, thus safeguarding the model’s integrity. Furthermore, Sanitizer has a negligible impact on the fidelity of the FL-trained model. Although there is a slight drop in ACC hovering around 5%, the trade-off is acceptable in our context in light of the substantial reduction in ASR.
Collectively, Sanitizer offers strong effectiveness in making watermarks harmless in FL and is widely applicable and not restricted to specific datasets.